# Overground gait transitions are not sharp but involve gradually changing walk-run mixtures even over long distances

### v2

@@ 1 Introduction
Humans and many other terrestrial animals exhibit a number of different gaits [1, 2, 3, 4, 5]. Humans walk, run, and much more occasionally, skip [6, 7]. Horses walk, trot, canter, and gallop, and more occasionally, use other gaits [3, 8, 9]. Transitions between such gaits have most commonly been studied using treadmills. In these treadmill gait transition experiments (Figure 1), the treadmill speed is changed slowly, either in a continuous fashion with some fixed acceleration, or in a series of acceleration phases alternating with constant speed phases [10, 11, 12, 13, 14, 15, 16, 17, 18]. These experiments found that people switch between walking and running around 2 m/s, but sometimes, the walk to run transition speed was different from — and higher than — the run to walk transition speed [18, 11, 12]. Importantly, the transition speeds were different from that predicted by energy optimality [10, 19], even though there is extensive evidence indicating energetics as a determinant of locomotor choice in numerous other settings [20, 21, 22, 23, 24, 25, 26, 2, 27, 23, 28, 29]. In such treadmill experiments, the gait transition is ‘sharp’, that is, happens at a particular speed where there is a preference of running over walking, or vice versa. Of course, humans do not spend their lives on treadmills, so their behavior may not already be energy optimal for such gait transition tasks without considerable learning and experience near the gait transition speeds [30, 31, 32]. Here, in contrast to these treadmill gait transition experiments, we show that overground gait transitions in realistic overground locomotion is more gradual and provide some clues for why there might exist distinct walk-to-run and run-to-walk speeds on a treadmill.
@@ 2 Methods
For each total distance $D_{\mathrm{total}}$, we used four different $T_{\mathrm{total}}$ corresponding to four different average speeds: 1.92, 2.24, 2.62, and 2.98  ms-1. These speeds were chosen to be roughly in the treadmill gait transition regime [5, 33, 10, 11], as well as the regime in which humans used a mixture of walking and running in our earlier study where we examined this protocol for a shorter distance [1]. Because of the large distances involved here, each participant either performed two trials both at the 2400 m distance, or three trials, two at 800 m and one at 2400 m. In all, we performed $N_{\mathrm{trial}}=$ 42 total trials across the 14 participants, with at least 9 trials per average speed. The trial assignment and trial order were randomized for each participant and across participants. The total duration was not revealed to the participants until a few moments before starting to eliminate as much pace planning as possible. All trials were performed on a 400 m outdoor track. Locomotion speeds were measured using GPS (10 Hz, VBOX Mini, Racelogic) and was available for all but one subject for whom this data was corrupted. Trials were video recorded and used to determine durations of walking and running during each trial, as visually classified by the investigators. The investigators classified the trial as running if there was a flight phase or if the hip moved downward and upward on a compliant-looking leg during stance phase, even if there did not seem to be a flight phase (jogging or grounded running); walking bouts are the complements of these running bouts [34, 35].
@@
These calculations are identical to that in [1] specialized to the task conditions herein, and are performed using a nonlinear optimization procedure. Here, in addition, noting that the energy landscape tends to have relatively flat minima [21, 41, 20, 31], we provide a sensitivity analysis: we compute the set of all speeds and gait durations that are within 1% and 2% of the previously computed optimal energy cost. To do this, we scan over a range of walking time durations for each speed, and solve for the optimal walking and running speeds for each such walking duration along with the corresponding energy costs, which then allows us to characterize the speeds and durations within 1% or 2% from the optimal energy cost.
@@
Figure 2: Humans use a walk-run mixture. a) Fraction of time spent running for the 800 m trials (blue dots) and the 2400 m trials (orange dots). Fraction of running increases with average speed. b) Number of switches from running to walking or walking to running. The median running fraction (green line) and 25-75th percentile (light green band) are also shown in panels a-b. c) Histogram of time spent at various speeds for trials at each of the four average speeds. The histograms pool data over all participants and all trials at the specific average speed. Each histogram has two peaks, one corresponding to walking (low speed) and another corresponding to running (high speed).
@@ 3 Results
Humans can achieve different average speeds with great precision over long distances.
Participants typically completed the trials in exactly the prescribed amount of time, arriving neither too early nor too late, using walking and/or running, appropriately adjusting their speeds as needed. Only in 3 trials out of 42 trials did the participants travelled so quickly initially that they had to be at virtually zero velocity for more than 10 sec, out of a typical trial duration of at least 540 seconds.
@@
Figure 3: a) Distribution of walking and running speeds changes gradually with average speed, with more time spent walking at low speeds and more time spent running at high speeds. Running speeds are systematically higher than walking speeds at each average speed. The overlap observed between walking and running speeds is largely an artifact of pooling participants (see Supplementary Figure S1). b) Gait transition speeds obtained from a logistic classifier do not change significantly speed. c) Mean running speeds increase with average speed. d) Mean walking speed remains largely constant or slowly decreases.
@@
The speed distributions during each trial have two peaks (Figure 2b). After categorizing every moment as either walking or running, we obtain separate walking speed and running speed histograms for each trial. These gait-specific histograms illustrate that in each of the walk-run mixtures, walking generally happens at low speeds and running generally happens at higher speeds (Figure 3a; see Supplementary Figure S1 for individual trial histograms). Specifically, people accomplish the required average speed by averaging over a slow walking speed and a faster running speed, as qualitatively predicted by energy optimality [1]. As the average speed increases, the speed distribution changes so that there is more time spent at the higher running speeds, resulting in a larger peak at the higher speed: speed 1.98 ms-1 is dominated by walking and speed 2.96 ms-1 is dominated by running.
@@
Computational predictions from energy optimality show that it is optimal to use walking alone for locomotion tasks with low average speeds (too much time), it is optimal to use running alone for tasks with high average speeds (very little time), and it is optimal to use a walk-run mixture for intermediate optimal speeds, as observed here (figure 4a,b). In the walk-run mixture regime, the model predicts that the walking fraction is high for low average speeds and the running fraction is high for high average speeds. Further, the model predicts that in the walk-run mixture regime, the walking speeds remains nearly constant but the running speed increases with average speed (figure 4a,b), as qualitatively observed in experiments (figure 3b-c). The running speed increasing with average speed requires there to be a cost for speed transitions [21], as not having any transition costs results in constant walking and running speeds across the different average speeds (see [1]). This result reemphasizes the importance of such transient costs in predicting human behavior.
Our sensitivity analysis reveals that the set of speeds and durations within 1% or 2% of the optimal energy costs is relatively broad, revealing a flat energy landscape, allowing humans high variability in speed with little energy penalty (figure 4b-c). This analysis shows that the energy cost is less sensitive to the running speed (larger band of speeds within 1% or 2% of optimality) compared to the walking speed; this lower sensitivity is inherited from the lower curvature of the running metabolic rate [5], so even normal running cost of transport is only weakly sensitive to running speed [39].
@@ 4 Discussion
We have performed overground gait transition experiments over much longer distances than earlier [1] and shown that humans use remarkably similar behavior to short distance bouts. Specifically, we find a gradually morphing walk-to-run transition regime, that involves walk-run mixtures, dominated by walking at lower speeds and by running at higher average speeds. Further, in contrast to treadmill gait transition experiments, here, humans do not use walking up to some instantaneous speed and running above that speed. Instead, within each trial, the overground walk-run mixtures involve transitions from walking at a low speed directly to running at a high speed and vice versa, generally with a gap between these speeds. The avoidance of a range of intermediate speeds has been previously found in free-living animals such as ostriches, emus, horses and gnus [42, 8, 43, 44]. We might speculate that the hysteretic gait-transition behavior on a treadmill [18, 11, 12] could be an extrapolation of the overground strategy with such speed gaps, deployed to an unfamiliar treadmill setting.
@@
Aside from instantaneous energetics [5, 8], gait transition from walking to running has been attributed to muscle force-velocity behavior [45], interlimb coordination variability [46], mechanical load or stress [47, 18], cardiovascular loads [48], and cognitive or perceptual factors such as visual flow [49, 50], all examined on a treadmill; see [51] for a review. There is mixed evidence on whether cognitive load influences the walk to run gait transition [52, 53]. However, none of these instantaneous factors has been used to explain why there may be hystereses between the walk-to-run and the run-to-walk speeds on treadmills, while Diedrich and Warren [18, 54] provided a multiple-energy-minima-based explanation of such hysteresis. In contrast, gait transition mechanisms that are dynamical systems and stability-based [55, 18, 56, 57, 58] cannot immediately explain the natural overground transitions observed here as a function of average speed. Specifically, it is unclear how any of these ‘instantaneous’ theories can explain the gradual walk-run-mixture-based gait transition regime behavior we have demonstrated in this manuscript. Predicting a walk-run mixture over a full bout as being optimal must necessarily require a theory that integrates some performance measure over the entire bout, rather than making decisions instantaneously based on crossing a threshold — or at the very least, use a control policy that switches gait and speed based on the time and distance remaining. Such control strategies could be explored in future work. Some recent work has also pointed to usefulness of time-like costs or discounting to explain gait-speed choices, in addition to energy, but given such considerations do not apply here due to the precise time durations provided for each locomotor task [59, 37].
@@
We have considered one kind of overground gait transition, in which the task is traveling a given distance at a desired average sub-maximal speed. Another kind of ecological overground gait transition that might have been common in our evolutionary past is to be walking normally and then having to accelerate to a higher running speed to either chase prey or evade a predator. Piers et al [60] considered such an overground gait-transition task, but theirs is a substantially different task from that considered in this manuscript: from an energy optimality perspective, the task of Piers et al [60] will be dominated by the cost of changing speeds and the cost of switching gaits [21, 38] — so we do not a priori expect identical gait transition speeds.
@@ Acknowledgment.
Data and code are available at this website for review purposes:
https://tinyurl.com/BakerWalkRunData and will be available publicly on Dryad upon publication.

### v3

@@ 1 Introduction
Humans and many other terrestrial animals exhibit a number of different gaits [1, 2, 3, 4, 5]. Humans walk, run, and much more occasionally, skip [6, 7]. Horses walk, trot, canter, and gallop, and more occasionally, use other gaits [3, 8, 9]. Transitions between such gaits have most commonly been studied using treadmills. In these treadmill gait transition experiments (Figure 1), the treadmill speed is changed slowly, either in a continuous fashion with some fixed acceleration, or in a series of acceleration phases alternating with constant speed phases [10, 11, 12, 13, 14, 15, 16, 17, 18]. These experiments found that people switch between walking and running around 2 m/s, but sometimes, the walk to run transition speed was different from — and higher than — the run to walk transition speed [18, 11, 12]. Importantly, the transition speeds were different from that predicted by energy optimality [10, 19], even though there is extensive evidence indicating energetics as a determinant of locomotor choice in numerous other settings [20, 21, 22, 23, 24, 25, 26, 2, 27, 23, 28, 29]. In such treadmill experiments, the gait transition is ‘sharp’, that is, happens at a particular speed where there is a preference of running over walking, or vice versa. Of course, humans do not spend their lives on treadmills, so their behavior may not already be energy optimal for such gait transition tasks without considerable learning and experience near the gait transition speeds [30, 31, 32]. Here, in contrast to these treadmill gait transition experiments, we show that overground gait transitions in realistic overground locomotion is more gradual and provide some potential clues for why there might exist distinct walk-to-run and run-to-walk speeds on a treadmill.
@@ 2 Methods
For each total distance $D_{\mathrm{total}}$, we used four different $T_{\mathrm{total}}$ corresponding to four different average speeds: 1.92, 2.24, 2.62, and 2.98  ms-1. These speeds were chosen to be roughly in the treadmill gait transition regime [5, 33, 10, 11], as well as the regime in which humans used a mixture of walking and running in our earlier study where we examined this protocol for a shorter distance [1]. Because of the large distances involved here, each participant either performed two trials both at the 2400 m distance, or three trials, two at 800 m and one at 2400 m. In all, we performed $N_{\mathrm{trial}}=$ 42 total trials across the 14 participants, with at least 9 trials per average speed and 10.5 trials per speed on average. The trial assignment and trial order were randomized for each participant and across participants. The total duration was not revealed to the participants until a few moments before starting to eliminate as much pace planning as possible. All trials were performed on a 400 m outdoor track. Locomotion speeds were measured using GPS (10 Hz, VBOX Mini, Racelogic, mounted in a waist pouch) and was available for all but one subject for whom this data was corrupted. Trials were video recorded and used to determine durations of walking and running during each trial, as visually classified by the investigators. The investigators classified the trial as running if there was a flight phase or if the hip moved downward and upward on a compliant-looking leg during stance phase, even if there did not seem to be a flight phase (jogging or grounded running); walking bouts are the complements of these running bouts [34, 35].
@@
These calculations are similar to that in [1] specialized to the task conditions herein, and are performed using the nonlinear optimization procedure fmincon in MATLAB; see Code availability for links to the MATLAB code. Here, in addition, noting that the energy landscape tends to have relatively flat minima [21, 41, 20, 31], we provide a sensitivity analysis: we compute the set of all speeds and gait durations that are within 1% and 2% of the previously computed optimal energy cost. To do this, we scan over a range of walking time durations for each speed, and solve for the optimal walking and running speeds for each such walking duration along with the corresponding energy costs, which then allows us to characterize the speeds and durations within 1% or 2% from the optimal energy cost.
@@
Figure 2: Humans use a walk-run mixture. a) Fraction of time spent running for the 800 m trials (blue dots) and the 2400 m trials (orange dots). Fraction of running increases with average speed. The median running fraction (green line) and 25-75th percentile (light green band) are also shown. The labels 2x and 3x refer to the number of multiple coincident data points, also indicated by sideways presentation jitter of the data points, though the prescribed averaged speed is just one of four possibilities indicated in Table 1. All 42 data points are represented here. b) Number of switches from running to walking or walking to running. The median number of switches (green line) and 25-75th percentile (light green band) are also shown. Labels 2x, 3x, etc., and sideways presentation jitter of data points added to indicate multiple coincident data points as in panel-a. Three trials in which participants did not reach on time are not plotted. c) Histogram of time spent at various speeds for trials at each of the four average speeds. The histograms pool data over all participants and all trials at the specific average speed. Each histogram has two peaks, one corresponding to walking (low speed) and another corresponding to running (high speed).
@@ 3 Results
Humans can achieve average speeds with precision using a clock over long distances.
Participants typically completed the trials in exactly the prescribed amount of time, arriving neither too early nor too late, using walking and/or running, appropriately adjusting their speeds as needed. Only in 3 trials out of 42 trials did the participants travelled so quickly initially that they had to be at virtually zero velocity for more than 10 sec, out of a typical trial duration of at least 540 seconds. This precision of arrival is perhaps to be expected given the countdown clock providing feedback on time remaining, the visible course providing feedback on the distance remaining, and the ability to change speed.
@@
Figure 3: a) Distribution of walking and running speeds changes gradually with average speed, with more time spent walking at low speeds and more time spent running at high speeds. Running speeds are systematically higher than walking speeds at each average speed. The overlap observed between walking and running speeds is largely an artifact of pooling participants (see Supplementary Figure S1). b) Gait transition speeds obtained from a logistic classifier do not change significantly speed. c) Mean running speeds increase with average speed. Mean walking speed remains largely constant or very slowly decreases.
@@
The speed distributions during each trial have two peaks (Figure 2c). After categorizing every moment as either walking or running, we obtain separate walking speed and running speed histograms for each trial. These gait-specific histograms illustrate that in each of the walk-run mixtures, walking generally happens at low speeds and running generally happens at higher speeds (Figure 3a; see Supplementary Figure S1 for individual trial histograms). Specifically, people accomplish the required average speed by averaging over a slow walking speed and a faster running speed, as qualitatively predicted by energy optimality [1]. As the average speed increases, the speed distribution changes so that there is more time spent at the higher running speeds, resulting in a larger peak at the higher speed: speed 1.98 ms-1 is dominated by walking and speed 2.96 ms-1 is dominated by running. While the pooled histograms show large gait-specific speed variability (Figure 3a), the individual-specific histograms have narrower speed distributions for each gait (Supplementary Figure S1).
@@
These qualitative features regarding walking speeds and running fractions are generally preserved across participants, even though there is considerable variability in the speeds and the running fractions between participants (Figure 2 and 3).
@@
Computational predictions from energy optimality show that it is optimal to use walking alone for locomotion tasks with low average speeds (too much time), it is optimal to use running alone for tasks with high average speeds (very little time), and it is optimal to use a walk-run mixture for intermediate optimal speeds, as observed here (figure 4a,b). In the walk-run mixture regime, the model predicts that the walking fraction is high for low average speeds and the running fraction is high for high average speeds. Further, the model predicts that in the walk-run mixture regime, the walking speeds remains nearly constant but the running speed increases with average speed (figure 4a,b), as qualitatively observed in experiments (figure 3b-c). The running speed increasing with average speed requires there to be a cost for speed transitions [21], as not having any transition costs results in constant walking and running speeds across the different average speeds (see [1] for the version without the transition costs). This result reemphasizes the importance of such transient costs [21] in predicting human behavior.
Our sensitivity analysis reveals that the set of speeds and durations within 1% or 2% of the optimal energy costs is relatively broad, revealing a flat energy landscape, allowing humans high variability in speed with little energy penalty (figure 4b-c). This analysis shows that the energy cost is less sensitive to the running speed (larger band of speeds within 1% or 2% of optimality) compared to the walking speed; this lower sensitivity is inherited from the lower curvature of the running metabolic rate [5], so even normal running cost of transport is only weakly sensitive to running speed [39]. Similarly, the near speed independence of running cost per unit distance (low curvature cost landscape) coupled with the curvilinear speed dependence of the walking costs may be responsible for the observed trends in the walking and running speed — that the running speed increases and the walking speed is relatively constant.
@@ 4 Discussion
We have performed overground gait transition experiments over much longer distances than earlier [1] and shown that humans use remarkably similar behavior to short distance bouts. Specifically, we find a gradually morphing walk-to-run transition regime, that involves walk-run mixtures, dominated by walking at lower speeds and by running at higher average speeds. Further, in contrast to treadmill gait transition experiments, here, humans do not use walking up to some instantaneous speed and running above that speed. Instead, within each trial, the overground walk-run mixtures involve transitions from walking at a low speed directly to running at a high speed and vice versa, generally with a gap between these speeds. The avoidance of a range of intermediate speeds has been previously found in free-living animals such as ostriches, emus, horses and gnus [42, 8, 43, 44]. We might speculate that the hysteretic gait-transition behavior on a treadmill [18, 11, 12] could be an extrapolation of the overground strategy with such speed gaps, deployed to an unfamiliar treadmill setting. That is, in our overground experiment, the walking speed when humans switch from walking to running is lower than the running speed when humans switch from running to walking, reminiscent of the hysteretic gait transition on a treadmill.
@@
Aside from instantaneous energetics [5, 8], gait transition from walking to running has been attributed to muscle force-velocity behavior [45], interlimb coordination variability [46], mechanical load or stress [47, 18], cardiovascular loads [48], and cognitive or perceptual factors such as visual flow [49, 50], all examined on a treadmill; see [51] for a review. There is mixed evidence on whether cognitive load influences the walk to run gait transition [52, 53]. However, none of these instantaneous factors has been used to explain why there may be hystereses between the walk-to-run and the run-to-walk speeds on treadmills, while Diedrich and Warren [18, 54] provided a multiple-energy-minima-based explanation of such hysteresis. In contrast, gait transition mechanisms that are dynamical systems and stability-based [55, 18, 56, 57, 58] cannot immediately explain the natural overground transitions observed here as a function of average speed. Specifically, it is unclear how any of these ‘instantaneous’ theories can explain the gradual walk-run-mixture-based gait transition regime behavior we have demonstrated in this manuscript. Predicting a walk-run mixture over a full bout as being optimal must necessarily require a theory that integrates some performance measure over the entire bout, rather than making decisions instantaneously based on crossing a threshold — or at the very least, use a control policy that switches gait and speed based on the time and distance remaining. Similarly, within each gait, there are speed fluctuations which may be partly due to feedback control based on the countdown clock and distance remaining, possibly correcting for any errors to arrive on time [59, 60, 61]. Such feedback control strategies and any learning effects over multiple trials could be explored in future work. It may also be fruitful to compare the structure of these speed fluctuations and control strategies to those during preferred walking overground walking without a time constraint [62], which does have higher instantaneous speed fluctuations than during walking on a treadmill [63].
Some recent work has also pointed to usefulness of time-like costs or discounting to explain gait-speed choices, in addition to energy, but given such considerations do not apply here due to the precise time durations provided for each locomotor task [64, 37]. One study [65] included time-related costs to explain the potential preference between walking and running based on distance traveled; in contrast to our study, this study examined fixed gait per bout (allowing no mixtures and encouraging constant speed) and examined explicitly ‘expressed preference’ (what people said they preferred based on trying both choices in a preference experiment) rather than ‘revealed preference’ (what people naturally do under the condition).
@@
We have considered one kind of overground gait transition, in which the task is traveling a given distance at a desired average sub-maximal speed. Another kind of ecological overground gait transition that might have been common in our evolutionary past is to be walking normally and then having to accelerate to a higher running speed to either chase prey or evade a predator. Piers et al [66] considered such an overground gait-transition task, but theirs is a substantially different task from that considered in this manuscript: from an energy optimality perspective, the task of Piers et al [66] will be dominated by the cost of changing speeds and the cost of switching gaits [21, 38] — so we do not a priori expect identical gait transition speeds between their and our experiments.
@@ Acknowledgment.
Data and code are available publicly at the Zenodo data repository:
https://doi.org/10.5281/zenodo.17103631